The height, in feet, of a rocket t secods after blast-off is given by a formula h=1260t-16t^2. After how many seconds will the rocket first reach a height of 23,400 feet?
Answer by checkley71(8403)

23,400=1260t-16t~2 or 16t~2-1260t+23,400=0 or (t-30)(t-48.75)=0 then
t=30 & t=4875 correct answe is 30 seconds.
proof 23,400=1260*30-16(30*30) or 23,400=37,800-16*900 or
23,400=37,800-14,400 or 23,400=23,400